An astronaut on a planet with no atmosphere drops a feather into a 5.90m deep crater and records that the feather falls freely for 1.64s.

What is the magnitude of free-fall acceleration on the planet?

With what speed does the feather strike the bottom of the crater?
1
Expert's answer
2020-12-14T07:27:05-0500

a) "h=gt^2/2\to g=2h/t^2=2\cdot5.9/1.64^2=4.39(m/s^2)"


b) "v=gt=4.39\cdot1.64=7.2(m/s)"